### v2

# Polaritonic Control of Blackbody Infrared Radiative Dissociation

Abstract
Vibrational strong light-matter coupling offers a promising approach for controlling chemical reactivity with infrared microcavities. While recent research has examined potential mechanisms for this phenomenon, many important questions remain, including what type of reactions can be modified and to what extent this modification can be achieved. In this study, we explore the dynamics of Blackbody Infrared Radiative Dissociation (BIRD) in microcavities under weak and strong light-matter interaction regimes. Using a Master equation approach, we simulate the effects of infrared field confinement and vibrational strong coupling on BIRD rates for diatomic molecules. We present a framework explaining how infrared microcavities influence BIRD kinetics, highlighting the importance of overtone transitions in the process. Our findings outline the conditions under which significant enhancement or mild suppression of BIRD rates can be achieved, offering insights into the practical limitations and new strategies for controlling chemistry within infrared resonators.
I   Introduction
Chemical reaction control is critical in applications ranging from industrial catalysis to medicine. Conventional methods for manipulating reaction rates involve varying temperature, pressure, or catalyst composition. An unconventional approach for controlling thermal reaction kinetics has been recently reported [1, 2, 3, 4] employing infrared (IR) microcavities [5] consisting of two moderate-quality mirrors separated by a distance $L_{C}$ of the order of IR wavelengths resonant with typical molecular vibrations [6]. When a collection of molecules with sufficiently large IR oscillator strength is embedded in a resonant microcavity, vibrational strong coupling occurs as signaled by the emergence of hybrid quasiparticles (vibrational polaritons) consisting of a superposition of molecular and electromagnetic modes of the confined device [7, 8, 9]. Recent theoretical research has probed potential mechanisms for the observed polariton effects on chemical reaction rates[10, 11, 12, 13, 14, 15, 16, 4, 17, 18, 19]. Still, several fundamental questions remain [20, 21], including what types of reactions can be controlled with an optical microcavity and to what extent this control can be exerted. In this study, we investigate gas-phase unimolecular dissociation activated by the absorption of microcavity blackbody infrared radiation [22] under weak and strong coupling with a suitable material.
Blackbody infrared radiative dissociation (BIRD) was proposed by Perrin in 1913 as a mechanism for gas phase unimolecular dissociation [23]. However, experimental observations in the 1920s ruled BIRD out in favor of the collisional activation mechanism [24, 25, 26]. Several decades later, the BIRD hypothesis was proven appropriate at sufficiently low pressures[27, 28, 29]. Since then, BIRD has become a valuable tool for investigating the thermal dissociation kinetics of molecular ions and clusters and has been observed in several simple and complex systems, including ion clusters, transition-metal complexes, and biopolymers [22].
The thermal radiation spectrum heavily influences BIRD rates. Given that free space and microcavity electromagnetic density of states and thermal radiation density can be significantly different [5], microcavity BIRD is expected to proceed with different rates compared to free space BIRD. However, the extent of this change, whether suppression or enhancement can be achieved, and if strong light-matter coupling introduces new possibilities relative to the weak coupling regime remain open questions. Our study provides answers to these questions. Our quantitative analysis focuses on the dissociation of diatomic molecules induced by multiphoton or multipolariton absorption in microcavities, but several qualitative conclusions are generalizable to polyatomic systems.
The primary aim of this study is to characterize thermal radiative dissociation in microcavities. We employ a Master equation treatment to investigate this process in (i) free space, (ii) under weak coupling with a nearly empty microcavity, and (iii) under weak coupling with a polaritonic material. Our results give upper bounds for BIRD enhancement and suppression in weakly and strongly coupled microcavities and identify overtone transitions as crucial ingredients for the observed modulation of BIRD. In Sec. II, we describe our methodology, Sec. III provides the main results and related discussion, and Sec. IV summarizes our main results and conclusions.
II   Methods
Figure 1: (a) Illustration of the Morse potential describing the vibrational states and transitions of the HF molecule. (b) and (c): A scheme illustrating an HF molecule inside a planar cavity and its dispersion under the weak coupling regime, respectively. (d) and (e): A scheme illustrating an HF molecule inside a planar cavity with the host material strongly coupled to the cavity and the dispersion of polariton states, respectively.
Master Equation for diatomic BIRD
Our analysis of BIRD in microcavities employs similar tools to those successfully applied to investigate this process in free space [28, 30, 31, 22]. These include a Markovian Master equation describing the time-evolution of vibrational quantum state populations with energy below the dissociation threshold defined below. Rotational degrees of freedom are ignored since they are weakly coupled to vibrational transitions, and the rotational dynamics are essentially classical at the temperatures of interest to BIRD. We also neglect Doppler and lifetime broadening throughout. The impact of nonzero vibrational linewidths on the presented results is described in Results and Discussion.
For the sake of simplicity, we assumed the thermalization of the background radiation is faster than all considered radiative transitions. The rate of change of the population in the $i$th vibrational state ($N_{i}$) is given by
$\frac{dN_{i}(t)}{dt}=\sum_{j\neq i}^{i_{\text{max}}}\left[k_{ij}N_{j}(t)-k_{ji}N_{i}(t)\right]-k_{\text{loss}}N_{i}(t)\delta_{i,i_{\text{max}}}\;,$ (1)
where $k_{ij}$ denotes rate of absorption when $i>j$ and emission when $i<j$ and $i_{\text{max}}$ is the highest energy bound state (Fig. 1a). This state is only loosely bound and is assumed to undergo irreversible dissociation at a rate $k_{\text{loss}}\gg k_{ij}\,\forall~{}i,j$. Equation 1 can be written in matrix notation as
$\frac{d[\textbf{N}]}{dt}=-\textbf{J}\cdot[\textbf{N}],$ (2)
where N is a column vector containing the vibrational populations and $\mathbf{J}$ is the transition matrix. The lowest eigenvalue $\lambda_{0}$ of J is the rate constant of the unimolecular reaction [32, 33] of the described mechanism. Note that this mechanism does not consider dissociation from intermediate vibrational levels ($\nu<i_{\text{max}}$). Thus, our results are expected to be valid for BIRD promoted by multiphoton dissociation through a particular doorway state, which is the highest-energy bound state in our case.
Radiative transition rates
Assuming the molecular system that undergoes BIRD is weakly coupled to the radiation field, the transition rates $k_{ij}$ can be obtained from perturbation theory (Fermi’s Golden Rule [34, 35]) and written in terms of Einstein coefficients in free space [36, 37] and their generalizations in a weakly coupled and polaritonic microcavity. Rotational dynamics is much faster than thermal infrared radiative transition, so we employ isotropically averaged radiative rates in the long-wavelength limit (electrical dipole approximation) throughout this work.
In free space, the spontaneous emission rate is given by the Einstein $A$ coefficient[36, 37]. In the case where spontaneous emission induces the $i\rightarrow j$ transition with $i>j$ this coefficient is given by
$k_{ij}^{\text{SpEm}}=A_{ij}=\frac{\omega_{ij}|\mu_{ij}|^{2}\pi}{3\epsilon_{0}\hbar}D_{0}(\omega_{ij}),$ (3)
where $\epsilon_{0}$ and $\hbar$ are the vacuum electrical permittivity and reduced Plank’s constant, respectively, $\omega_{ij}=(E_{i}-E_{j})/\hbar$ is the transition frequency, $\mu_{ij}$ is the corresponding transition dipole moment and $D_{0}(\omega)=\omega^{2}/\pi^{2}c^{3}$ is the free-space radiative density of states (DOS). The Einstein $B$ coefficient associated with stimulated emission and absorption processes involving states $i$ and $j$ is given by $B_{ij}=\pi|\mu_{ij}|^{2}/(3\epsilon_{0}\hbar^{2})$. The corresponding thermal stimulated emission and absorption rate constants are given by
$k_{ij}^{\text{Abs}}(T)=k_{ij}^{\text{StEm}}(T)=B_{ij}\rho_{0}(\omega_{ij},T)\;,$ (4)
where $\rho_{0}(\omega,T)$ is the free space blackbody radiation energy density
$\rho_{0}(\omega,T)=\hbar\omega~{}n_{\text{BE}}(\omega,T)D_{0}(\omega),$ (5)
and $n_{\text{BE}}(\omega,T)$ is the Bose-Einstein distribution thermal occupation number $n_{\text{BE}}(\omega,T)=1/[\text{exp}(\hbar\omega/k_{B}T)-1]$[38] and $k_{B}$ is the Boltzmann constant.
The final pieces necessary to compute the rate constants above are the transition dipole matrix $\mu_{ij}$ and the frequencies $\omega_{ij}$. We estimate the latter assuming the electronic ground-state potential energy curve of the examined diatomic molecules is given by the Morse potential [Fig. 1(a)][39]. The bound wave functions associated with the Morse potential are also employed in the computation of the vibrational transition dipole moments $\mu_{ij}$ via
$\mu_{ij}$ $=\int\psi^{*}_{i}(r)\mu(r)\psi_{j}(r)\mathrm{d}r,$ (6)
where $\psi_{i}(r)$ represents the $i$th vibrational level wave function, and the dipole function $\mu(r)$ models the nonlinear change in the (electronic ground-state) molecular electrical dipole with the bond length $r$ [40].
We investigated the IR radiative dissociation of hydrogen fluoride, HF, and lithium hydride, LiH. The employed dipole function $\mu(r)$ for HF is $\mu(r)=\mu_{0}re^{-\zeta r^{4}}$, where $\mu_{0}=0.45414e$ and $\zeta=0.081616$-4 were obtained from previous work on laser-induced HF dissociation[41]. The LiH dipole function was interpolated from ab initio quantum chemical computations (CCSD(T)/aug-cc-pVTZ, see Supplementary Information for further details). Morse potential parameters include the equilibrium bond length ($r_{e}$), electronic ground-state dissociation energy ($D_{e}$), and harmonic frequency ($\omega_{e}$) of each molecule and are listed in Table 1.
Table 1: Parameters used for Morse potential of HF [42, 43, 44, 41] and LiH [45].
Constant HF LiH
$D_{e}$ (eV) 6.123 2.641
$r_{e}$ (Å) 0.926 1.595
$\nu_{e}$ (cm-1) 4138 1405
$\nu_{e}\chi_{e}$ (cm-1) 86.7 28.3
$m_{r}$ (a.u) 0.95 0.88
The examined dissociation processes (via the most energetic bound state) have reduced room temperature rates relative to typical BIRD experiments involving polyatomic ions [27, 46, 29, 22] due to the relatively high stability and restricted number of state-space pathways available for diatomic dissociation proceeding via the highest energy level. This feature motivated our choice of examining diatomic BIRD at temperatures $T=\mathcal{O}(\hbar\omega_{e}/k_{\text{B}})$, an order of magnitude greater than room temperature. The SI shows that the trends discussed below hold for a wide interval of temperatures. Despite their different energetics, microcavity effects on HF and LiH BIRD yield the same qualitative behavior.
The transition rates $k_{ij}$ are proportional to the electromagnetic density of states $D(\omega)$, suggesting BIRD rates can be controlled by placing the reactive molecule inside a microcavity. To examine this scenario, we adapt the free space formalism described above to arbitrary photonic devices by replacing $D_{0}(\omega)$ in Eqs. 3 and 5 with the appropriate DOS for a weakly coupled microcavity $D_{\text{C}}(\omega)$.
In a planar empty Fabry-Perot microcavity composed of two parallel mirrors (with unit reflectivity for simplicity) separated by a distance $L_{C}$, [Fig. 1(b)], the frequencies of the electromagnetic modes are given by
$\omega_{\text{C}}(m,q)=c\sqrt{q^{2}+\frac{m^{2}\pi^{2}}{L_{C}^{2}}}\;,$ (7)
where $\mathbf{q}=(q_{x},q_{y})\in\mathbb{R}^{2}$ is the in-plane projection of the mode wave vector, $q^{2}=q_{x}^{2}+q_{y}^{2}$, and $m$ is an integer quantum number that identifies the longitudinal wave vector $k_{z}=m\pi/L_{C}$ and thus the photonic energy bands in Fig. 1(c). For transverse-magnetic (TM) modes $m=0,1,2,...$ whereas $m=1,2,3...$ for transverse-electric (TE) modes [47, 48]. Including the TM0 mode here is crucial, as otherwise, transitions with energy below the cavity cutoff would be impossible.
Using Eq. 7, it can be shown (see Supporting Information) that the spatially-averaged electromagnetic DOS inside a microcavity in the weak coupling regime can be expressed by [49]
$D_{C}(\omega)=\frac{\omega}{\pi c^{2}L_{C}}\left\lfloor\frac{\omega L_{C}}{\pi c}\right\rfloor+\frac{\omega}{2\pi c^{2}L_{C}}\;,$ (8)
where $\lfloor x\rfloor$ denotes the floor function.
Lastly, we considered the BIRD scenario in which the diatomic interacts weakly with the electromagnetic component of polariton modes emergent from vibrational strong coupling between the microcavity modes and a material system (solid-state or molecular ensemble) with sufficiently large IR oscillator strength. We emphasize this material system does not undergo BIRD; its role is to generate polariton normal-modes with a tunable density of states and thermal energy density, as shown in Figs. 1(d) and (e). The formalism discussed above for BIRD in a weakly coupled microcavity can be straightforwardly adapted to model polariton-mediated BIRD (see SI) by replacing the photon density of states $D_{0}(\omega)$ with the photon-weighted polariton density of states $D_{P}(\omega)$.
The Hamiltonian of the strongly coupled molecular subsystem and the microcavity is given in the Power-Zienau-Woolley (dipole) gauge [50, 51]
$H=H_{\text{SM}}+H_{\text{L}}+H_{\text{INT}}\;,$ (9)
where the interaction $H_{\text{INT}}$ is given by
$H_{\text{INT}}=\frac{1}{\epsilon_{0}}\int\left[-\textbf{D}(\textbf{r})\cdot\textbf{P}(\textbf{r})+\frac{1}{2}\textbf{P}^{2}(\textbf{r})\right]d^{3}\textbf{r},$ (10)
[…]
$D_{P}(\omega)=\sum_{m=0}^{\infty}\left(1-\frac{\delta_{0,m}}{2}\right)\frac{q(m,\omega)P_{C}(\omega)}{\pi v_{g}(\omega,q)L_{C}}\theta[\omega_{\text{C}}(\omega)-\frac{m\pi c}{L_{C}}],$ (11)
where $\theta(x)=1$ for $x\geq 0$ and vanishes elsewhere, $\omega_{\text{C}}(\omega)$ is the frequency of the photon mode that under strong coupling leads to the formation of a polariton with frequency $\omega$, $P_{C}(\omega)$ is the frequency-dependent photon content of polaritons with frequency $\omega$
$P_{C}(\omega)=\frac{(\omega^{2}-\omega_{M}^{2}-\Omega_{R}^{2})^{2}}{(\omega^{2}-\omega_{M}^{2}-\Omega_{R}^{2})^{2}+\omega^{2}_{C}(\omega)\Omega_{R}^{2}},$ (12)
$\omega_{M}$ is the matter oscillator frequency, $\Omega_{R}$ is the collective light-matter interaction strength, and the polariton group velocity ($v_{g}$) is given by
$v_{g}(\omega,q)=\frac{qc^{2}}{2\omega}\left(1\pm\frac{\omega_{C}^{2}-\omega_{M}^{2}+\Omega_{R}^{2}}{\sqrt{\left(\omega_{C}^{2}+\omega_{M}^{2}+\Omega_{R}^{2}\right)^{2}-4\omega_{C}^{2}\omega_{M}^{2}}}\right)\;,$ (13)
where the positive sign applies when $\omega$ belongs to an upper polariton branch, i.e., $\omega>\omega_{C}$, and the negative sign is used if $\omega<\omega_{C}$. In this expression, we have omitted the frequency dependence of $\omega_{C}$ for simplicity.
Our model shows singular behavior for $D_{P}(\omega)$ as $\omega\rightarrow\omega_{M}$. In this limit, both the photon content $P_{C}(\omega)$ and the group velocity $v_{g}(\omega,q)$ approach zero, but the ratio $P_{C}(\omega)/v_{g}(\omega,q)\rightarrow\infty$ resulting in $D_{P}(\omega)$ diverging to positive infinity. This singularity is due to the formation of the well-known stopgap region[59, 60, 61, 62] where $D_{P}(\omega)=0$. As previous work shows [63], introducing an ultraviolet cutoff and material disorder eliminates the singularity of $D_{P}(\omega)$ at $\omega_{M}$. Nevertheless, there remains significant enhancement of $D_{P}(\omega)$ in a neighborhood of $\omega_{M}$, which depends in a non-universal way on the type of disorder and coherence loss mechanism. This motivates our decision to use the formally exact expression presented in Eq. 11 and interpret our results as upper bounds on the effect of polaritons on BIRD.
III   Results and Discussion
Microcavity-assisted blackbody infrared radiative dissociation
Our analysis of radiative dissociation in weakly coupled microcavities assumes a single diatomic molecule (or a dilute sample at low enough pressure that collisions can be neglected) in a planar microcavity. The tunable electromagnetic environment afforded by the microcavity allows BIRD rates to be either enhanced or suppressed relative to free space, depending on which transitions are significantly affected by the modified photon DOS.
In Fig. 2, BIRD rates in the weak coupling regime are shown relative to the free space rate ($k_{0}=2.70\times 10^{-6}$ s-1 for HF at $T/\omega_{e}=0.67$ corresponding to $T=4000~{}\text{K}$ and $k_{0}=1.64\times 10^{-5}$ s-1 for LiH at $T/\omega_{e}=0.99$ corresponding to $T=2000~{}\text{K}$). For small microcavity lengths, i.e., $L_{C}<2$ $\mu$m for HF and $L_{C}<5$ $\mu$m for LiH, a mild but consistent enhancement of BIRD rates is observed. At these lengths, an overall increase in the microcavity electromagnetic DOS over a wide frequency range covering several important transitions is associated with the observed enhancement. The ratios of microcavity BIRD rates to free space BIRD rates quickly approach 1.0 as $L_{C}$ increases beyond a few microns. This is expected since the microcavity DOS converges to the free space limit when $L_{C}\rightarrow\infty$.
The complex oscillatory pattern in the BIRD rates relative to free space shown in Fig. 2 arises from the interplay of discrete vibrational transitions enhanced or suppressed depending on $L_{C}$. As we demonstrate below, this complexity is caused by overtones, which play an essential role in our analysis.
Figure 2: Ratio of BIRD rate inside a microcavity (weak coupling regime) to the free-space rate as a function of microcavity length for HF and LiH molecules. Red circles highlight specific lengths further discussed in Fig. 3.
We conducted a sensitivity analysis of the dissociation rate to interpret the origin of the oscillatory behavior in Fig. 2 and identify the most consequential radiative transitions involved in the BIRD of HF and LiH. This analysis measures how impactful an enhancement in the photon DOS at a specific transition frequency is to the dissociation process (see Supporting Information for more details). For HF dissociation, changes in the photon density of states corresponding to the frequency of overtone transitions $i\rightarrow 23$, with $i\in\{16,17,19\}$ (in the relevant 1500 -- 5000 cm-1 frequency range) are the most consequential for the BIRD rate. Similarly, for LiH, the relevant transitions are $i\rightarrow 29$, with $i\in\{21,22,23,25\}$. In both cases, the identified overtones take the vibrational mode to the highest energy-bound state (23 for HF and 29 for LiH), from which dissociation rapidly occurs. We emphasize that overtones are crucial because, at high excitation levels, fundamental transitions ($i\rightarrow i+1$) involve very small energies and, therefore, have associated small thermal photon populations and reduced oscillator strengths. See Supporting Information for a brief discussion of results obtained without overtone transitions.
Figure 3: Microcavity photon DOS normalized to the free space DOS for selected lengths ($L_{C}$) where BIRD suppression and enhancement were observed for HF (left panels) and LiH (right panels) in Fig. 2. The dotted vertical lines indicate the location of the transition frequencies corresponding to the most relevant overtone transitions, with color coding to distinguish between suppression (red) and enhancement (green). Panel (a) shows that the small HF BIRD suppression observed at $L_{C}=3.03~{}\mu\text{m}$ is at least in part caused by the reduced relative DOS at the highlighted overtones. Panel (b) shows the mild enhancement of HF BIRD in a microcavity with $L_{C}=3.21~{}\mu\text{m}$ can be ascribed to the greater microcavity photon DOS at the frequencies of the highlighted overtones. Panels (c) and (d) show the analogous scenario for LiH, where BIRD is observed to be suppressed at $L_{C}=9.4~{}\mu\text{m}$ but enhanced at $L_{C}=9.6~{}\mu\text{m}$, primarily due to the variation of the microcavity photon density of states at the selected $i\rightarrow 29$ overtone frequencies.
Equipped with the knowledge of the most relevant transitions for the investigated BIRD of HF and LiH via the most energetic bound state, we can readily explain the microscopic origin of the intricate pattern in Fig. 2 by considering how the microcavity DOS changes around the previously mentioned overtones. Fig. 3 presents the ratio of the microcavity DOS to the free space for selected values of $L_{C}$ where BIRD suppression or enhancement was verified in Fig. 2 (red circles). Overtone frequencies are highlighted with dotted vertical lines, color-coded by whether they are enhanced (green) or suppressed (red) in the corresponding microcavity. In Fig. 3(a) ($L_{C}=3.03$ $\mu$m), the most relevant HF overtones are suppressed, and this leads to the observed local minimum in $k_{c}/k_{0}$ in Fig. 2. In contrast, when the microcavity length slightly increases to $L_{C}=3.21$ $\mu$m [Fig. 3(b)] the photon DOS is enhanced at the $16\rightarrow 23$ and $19\rightarrow 23$ transitions frequencies, and we observe a corresponding leap in the BIRD rate in Fig. 2. Similarly, in Fig. 3(c) ($L_{C}=9.4~{}\mu\text{m})$, all of the relevant LiH overtones but $21\rightarrow 29$ are suppressed, leading to a reduced BIRD rate in this microcavity relative to free space, whereas Fig. 3(d) shows that at $L_{C}=9.6~{}\mu\text{m}$ most of the relevant LiH overtones are enhanced, thus explaining the sudden increase in BIRD rate at $L_{C}=9.6~{}\mu\text{m}$ relative to $L_{C}=9.4~{}\mu\text{m}$. Introducing a finite linewidth to the considered radiative transitions would smooth out the oscillations observed in Fig. 2 and dampen their amplitude. Consequently, the weakly coupled microcavity effect on BIRD would be significant only at lengths less than a few microns.
Note that suppression effects are generally minor in Fig. 2 because the dissociative process has many pathways. Therefore, blocking or slowing down a particular transition may reduce the dissociation rate, but the effect is almost always negligible as the molecular system explores alternative dissociation pathways. Our analysis of BIRD in weak coupling reveals two main results we expect to hold generically for chemistry in microcavities: (i) overtone transitions are crucial and should not be ignored, and (ii) blocking a single reactive pathway is likely inconsequential.
Polariton-assisted blackbody infrared radiative dissociation
To analyze polariton-assisted BIRD, we considered a microcavity containing, in addition to the reactive diatomic molecule, a host material strongly coupled with the resonator modes. Fig. 1(d) illustrates this scenario, emphasizing the formation of polariton states, which effectively change the photon frequency-momentum dispersion relation [Fig. 1 (e)] and the electromagnetic DOS. We explore the implications for diatomic BIRD via the most energetic bound state below.
Figure 4: Ratio of polariton-assisted BIRD rates to free space BIRD rates. In the polaritonic case, the diatomic molecule is embedded in a strongly coupled microcavity with variable host material with frequency $\omega_{M}$. The top panels, (a) and (b), correspond to HF, and the bottom panels, (c) and (d), show results for LiH. In (a) and (c), the collective light-matter interaction strength, $\Omega_{R}$, is fixed at 400 and 200 cm-1, respectively. The horizontal dashed gray and green lines indicate where the dissociation rates equal $k_{0}$ and $k_{c}$, respectively. The insets in (a) and (c) show zoomed-in views around the overtones $17\rightarrow 23$ for HF and $22\rightarrow 29$ for LiH. In (b) and (d), relative BIRD rates are shown for different Rabi frequencies ($\Omega_{R}$), depicting the dependency on the detuning between the host molecule and the diatomic transition energies.
Fig. 4 presents the ratio of polariton-assisted BIRD to the corresponding free space rates ($k_{p}/k_{0}$) as a function of the host matter frequency, $\omega_{M}$. These computations were performed with $\Omega_{R}=400~{}\text{cm}^{-1}$ and 200 cm-1 by fixing $L_{C}$ at 3.21 and 9.60 $\mu\text{m}$ for HF and LiH respectively. As discussed in the previous section, these choices for $L_{C}$ enhance the dissociation rate in the weak coupling regime. We observe prominent peaks for HF and LiH where the relative dissociation rate is notably enhanced. Maximal enhancement occurs when the matter frequency ($\omega_{M}$) is greater than but sufficiently close to a particular diatomic overtone transition frequency we denote by $\omega_{i\rightarrow j}$. Here, we focus on the case where $i=17$ and $j=23$ for HF and $i=22$ and $j=29$ for LiH. In contrast, the inset of Figs. 4(a) and (c) show that when $\omega_{M}$ approaches the diatomic transition from the left, the dissociation reaction is suppressed. This suppression occurs because as $\omega_{M}-\omega_{i\rightarrow j}\rightarrow 0^{-}$, the overtone transition falls into the stopgap region of the polaritonic dispersion, where the electromagnetic DOS is zero (Fig. 1e). However, the BIRD suppression is relatively small since, as discussed earlier, the reaction may still proceed through alternative pathways.
The effect of the collective coupling strength ($\Omega_{R}$) on the relative polariton-assisted BIRD rates is examined in Fig. 4(b) and (d). When $\omega_{M}-\omega_{i\rightarrow j}$ is around 0.01-1 cm-1, increasing $\Omega_{R}$ results in greater polariton-assisted BIRD rates. However, as $\omega_{M}$ becomes arbitrarily close to $\omega_{i\rightarrow j}$, the polaritonic enhancement of BIRD saturates around 20 and 50 for LiH and HF, respectively. This plateau appears to be insensitive to the magnitude of the Rabi splitting. It can be explained by the singularity in the photon-weighted polariton DOS, which arises as $\omega\rightarrow\omega_{M}$ and is illustrated in Fig. 5.
Figure 5: Photon-weighted polariton density of states $D_{P}(\omega)$, normalized to the free space DOS, for a microcavity with length $L_{C}=9.6~{}\mu$m strongly coupled to a material with a bright transition at frequency $\omega_{M}=1401$ cm-1. The curves show the effect of different Rabi splitting values ($\Omega_{R}$) on the photon-weighted polariton DOS, which is relevant for radiative processes mediated by polaritonic systems. As $\Omega_{R}$ increases, the frequency range below $\omega_{M}$, where the photon-weighted polariton DOS significantly exceeds that of free space, widens. This trend explains the impact of the collective light-matter interaction strength on the observed polariton-assisted bond infrared dissociation (BIRD) enhancement shown in Fig. 4.
In Fig. 5, the photon-weighted polariton density of states $D_{P}(\omega)$ is shown for selected values of $\Omega_{R}$ at $L_{C}=9.6$ $\mu$m and $\omega_{M}=v_{22\rightarrow 29}(\text{LiH})=1401$ cm-1. The case where $\Omega_{R}=0$ represents the weak coupling limit and is included here to directly compare $D_{P}(\omega)$ with $D_{C}(\omega)$. It can be seen from Fig. 5 that the photon-weighted polariton DOS is only fundamentally different from the bare microcavity DOS in a neighborhood of $\omega_{M}$ that also contains the stopgap. Most remaining features, e.g., the oscillations in $D_{P}(\omega)$ with $\omega$, are also observed in weak coupling and can be understood from Eq. 8. The various curves in Fig. 5 illustrate how the width of the stopgap region and the electromagnetic mode frequencies with significantly enhanced $D_{P}(\omega)$ expand as $\Omega_{R}$ increases. As a result, strong coupling can induce changes in the rates of multiple vibrational transitions depending on the Rabi splitting.
The singularity in the photon-weighted polariton DOS enables a particular radiative overtone transition to become arbitrarily faster than the corresponding free space rate. However, our results demonstrate that this singularity does not lead to arbitrarily fast reaction rates since the complex multi-step reactive process ensures that other transitions quickly become rate-limiting. Therefore, the BIRD rates reported as $\omega_{M}\rightarrow\omega_{i\rightarrow j}$ provide a finite upper bound for polariton-enhanced BIRD rates in the lossless limit discussed here. In experiments, imperfections in the electromagnetic device and the strongly coupled material smooth out the photon-weighted DOS near the stopgap. Nonetheless, previously reported simulations show the enhancement of $D_{P}(\omega)$ persists around $\omega_{M}$ in the presence of an ultraviolet cutoff and material disorder [63]. The introduction of a UV cutoff, loss, and dephasing pathways for the material system will lead to reduced BIRD enhancements relative to those in Fig. 4 depending in a non-universal fashion on $\Omega_{R}$ and parameters characterizing the quality of the strongly coupled subsystems (e.g., microcavity quality factor, matter homogeneous linewidth, etc).
Including a finite linewidth for the diatomic vibrational transitions relaxes the strict resonance requirement discussed above ($\omega_{M}\rightarrow\omega_{i\rightarrow j}$), and the overall enhancement will depend on the spectral overlap between $\omega_{i\rightarrow j}$ and $D_{P}(\omega)$. Since the molecular linewidth and cavity imperfections are expected to be limiting factors, it is clear that our reported results represent theoretical upper bonds for polariton-assisted BIRD rates for the reaction mechanism investigated here.
Before ending, we emphasize our observation of the same qualitative behavior for polariton-mediated BIRD in LiH and HF. Since these chemical species have significantly different energetics, we anticipate the main trends reported in this work also apply to other diatomic and polyatomic molecules, particularly in cases where dissociation occurs through a specific doorway state.
IV   Conclusions
This work presented a detailed analysis of BIRD in infrared microcavities. Our analysis revealed mechanisms for controlling dissociation phenomena in microcavities, particularly by manipulating the electromagnetic density of states (DOS) in weak and strong light-matter coupling regimes. Overtones play a critical role in the dissociation process under both weak and strong coupling cases, as the corresponding transitions act as rate-limiting steps for the dissociative process occurring via the highest energy-bound state. The radiative dissociation rate changes induced by a microcavity are essentially tied to how the photonic density of states is modified around specific overtone transitions driving the system into the dissociating (doorway) state.
We reported microcavity-assisted dissociation rate enhancements of $O(1)$ in the weak coupling regime. In microcavities where $L_{C}\leq 5~{}\mu\text{m}$, the photon DOS is greater than in free space for most relevant diatomic transitions, leading to a mild increase in dissociation rates. As $L_{C}$ increases beyond a few microns, a complex pattern emerges from the interplay between enhanced and suppressed vibrational transitions, which can be rationalized by analyzing a few important overtones. Further increases in $L_{C}$ causes the BIRD rates to approach the free-space limit.
In the strong coupling regime, where polaritons mediate BIRD, we reported $O(10)$ enhancements when the host matter frequency is resonant with rate-limiting overtones of the dissociative process. This enhancement is due to the emergence of a singularity in the photon-weighted polariton DOS, which massively increases transition rates between levels with energy differences that approach the low-energy boundary of the polariton stopgap from below. The singularity is attributed to the lack of a high-energy cutoff for the electromagnetic field and the lossless character of the strongly coupled system. Nevertheless, while the singular behavior of the photon-weighted polariton DOS is absent from experiments, we have argued its effects are expected to persist to an extent that depends on $\Omega_{R}$ and parameters controlling losses and disorder. Ultimately, the BIRD enhancement factors reported here provide upper bounds for polariton-induced diatomic thermal radiative dissociation changes via the most energetic bound state.
Our work focused on how optical microcavities can modify BIRD reactions of diatomic molecules via a particular mechanism. Despite focusing on these specific systems, the main qualitative conclusions we drew about BIRD in weak and strong coupling apply broadly. We expect the analysis introduced here will be useful in understanding and controlling generic polyatomic thermal radiative-induced dissociation processes in microcavities.
Acknowledgments. This work was supported by the donors of ACS Petroleum Research Fund under Doctoral New Investigator Grant 66399-DNI6. R.F.R. served as Principal Investigator on ACS PRF-66399-DNI6 that provided support for E. S. R.F.R also acknowledges support from NSF CAREER award Grant No. CHE-2340746, and startup funds provided by Emory University.
Author Contributions. G. J. R. A. and E. S. contributed equally to this work.